Telescope and AGw4 Commissioning at LDG - 09-10 May 2009 MST

Observers:: J. Hill (mtn), A. Rakich (Tucson)
Telescope Operator: W. Wack
Software Support: N. Cushing (phone) Telescope Support: J. Little

Summary

Sky is photometric again, and seeing is good until the last hour of the night.

Gave up the first hour to shutter door alignment checks in return for having IRTC completely mounted and cabled at LDG focus.

Lost an hour to some missing GCS configuration files in Build 26. This was tangled up with some problems getting oacserver to properly transition from AGW3 to AGW4.

Andrew took extensive field aberration data on BS9152 with IDL. He was raving (in a happy way) about the quality/stability of the data while taking it. Then he took a smaller set of data with GCS. The preliminary result is that both IDL and GCS are now handling the LDG - AGW4 active optics data correctly. Doug and Andrew will analyze this data on Sunday in their respective daytimes.

While Andrew was taking data, John completed the transtar.cl script that collects AGW transformation data automatically. Got 16 good transform data points on StoneM before the seeing fell apart, and then a bunch more points in 1 arcsec seeing.

Preparations

Moved SX M3 swing arm to retracted position. Move IRTC2 to the LDG focus on AGW4.

John L., Jared and Rick working until 20:30 inspecting Left Shutter Door rails. This was a planned activity that was delayed an hour by the moving and re-cabling of IRTC. I didn't really mind, because the Penguins-Capitals game went into overtime. Penguins won! and now lead the series 3-2.

19:00 Rebooted agw-control to get rid of a "defunct" oacserver.

Details

Problems with GCS and AGW4

20:30 - Ready to observe, but GCSL crashes everytime we say 'selectAGW AIP_IRTC_DG_L.cfg'. We switched back to AGW3, and could select it OK. Then we had alot of problems getting back to AGW4. I have an impression that you have to stop oacserver when you switch AGW units, but I can't prove it. Then we were getting RPC errors on 'startAGW -u 1'. Were these from the MOXA??? Called Norm, and we got AGW4 running after another round of stopAGW / startAGW. Norm ran the debugger on the core file, and we figured out that the AIP_IRTC_DG_L_WFScam.ini and AIP_IRTC_DG_L_guidecam.ini CCD configuration files were missing from B26. Lost 1 hour to this problem.

Observing

21:30 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - forgot to authorize in IRAF

authorize DGTEST

21:31 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - Adjust IE=-44 CA=+4

21:37 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - got some M2 offload that shifted pointing 2 arcsec.

21:39 pointto BS9152 Tel="ACQUIRE" PAM- gs=0 - Adjust IE=-37 CA=+4

Active Optics verification (comparison between IDL and GCS)

21:43 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=2

21:50 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=18

21:56 pointto BS9152 Tel="ACTIVE" PAM+ PA=INDEF=338.61 gs=4

21:59 stop WFS in GCS, so we can take data with IDL (but there was a problem)

22:05 Restart WFS in GCS. Converges to 0.6 arcsec and 116 nm rms

22:11 pointto BS9152 Tel="ACTIVE" PAM+ PA=0

22:18 pointto BS9152 Tel="GUIDE" PAM+ PA=INDEF65.44 gs=18

22:19 Andrew collimating with IDL.

22:21 pointto BS9152 Tel="GUIDE" PAM+ PA=0 gs=0 - The Z6 measurement seems inconsistent.

22:26 pointto BS9152 Tel="ACQUIRE" PAM- PA=0 gs=0 - Adjust IE=-49 CA=+4

22:28 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=0 - Seeing is varying 0.5 to 0.6 arcsec.

22:32 Changed from 'split' mode to 'primary' mode. Sent focus 0 to force offload.

22:38 Send Z9=+300 nm. wfsc 0047

Field Aberration Measurements
Using IDL

BS9152
GS PA WFSC
0 0 21-22
17 124.77 23-25
0 124.77 26-29
18 65.44 30-32
0 65.44 33-35
7 10.54 36-38
0 10.54 39
4 338.61 40-42
0 338.61 43
1 229.72 44-46
0 229.72 47-54
9 168.12 55-57
0 168.12 58
17 124.77 59-61
0 124.77 62-69
18 65.44 70-72
0 65.44 73-74
7 10.54 75-77
0 10.54 78
4 338.61 79-81
0 338.61 82
1 229.72 83-85
0 229.72 86-87
9 168.12 88-90
0 168.12 91-92
17 124.77 93-95
0 124.77 96-103
18 65.44 104-105
0 65.44 106-111
7 10.54 112-114
0 10.54 115
4 338.61 116-118
0 338.61 119
1 229.72 120-122
0 229.72 123-125
9 168.12 126-128
0 168.12 129

Note only 2 measurements at star 18 third time around, operator error

Data lives in /Repository/AGW_Data/090510w/

AGW Transform Script Tests

01:18 John is back at the helm

01:20 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=8 verb- (gs 8 is a binary)

01:28 first test of transtar.cl for automated collection of AGW transformation data - accidentally guided in PARALLACTIC mode. It worked.

1:35 now doing a check of field aberration measurements collimating and measuring with GCS (and watching IDL sends compared to GCS sends)

Field Aberration Measurements
Using GCS (with varying exposure times from 15s-30s )

BS9152
GS PA WFSC
0 0 77-82
17 124.77 83-85
0 124.77 93-102
18 65.44 103-107
0 65.44 109-112
7 10.54 113-117
0 10.54 119-126
4 338.61 127-129
0 338.61 134-140
1 229.72 141-144
0 229.72 145-147
9 168.12 148-152
0 168.12 153-154

AGW4 - IRTC2 focus measurement

02:50 pointto StoneMagw Tel="ACQUIRE" PAM- gs=0 - Adjust IE=-48 CA=+10

02:56 pointto StoneMagw Tel="ACTIVE" PAM+ PA=0 gs=51

03:00 cl < focus_sequence.cl - AGW focus is 28.119 at star 51 corresponding to 31 on-axis. Focus was near the beginning of this sequence, and stars were close to saturation near focus. No need to reduce this dataset. Will repeat this tomorrow night after we collect the preliminary transform data for the rest of tonight.

NOTE: that LUCIFER focus is not up-to-date in AIP_L.cfg in Build 26, and presumably also Build 25. We need to trace back this history to the last measurement. JMH: I was wrong about this. 31.0 is the correct focus value for LUCIFER and AGW3 measured in mid-January 2009. Of course, it will need to be remeasured in warm weather.

AGW Transform Data

Starting a fresh version of /home/agwuser/logbook.1 tonight.

Data being written to /Repository/ but will wind up in /Repository/IRTC_Data/090510/

These data were taken with LEFTZEROPOINT = +5.5 degrees

03:10 transtar 40

This is the new automated script to collect transformation data. See the output below. What I haven't figured out yet is how to collect the log information in the wiki. Note that transtar puts the star number in the getdat results in logbook.1, and it always remembers to do getdat.

transtar presetting telescope to guide star 40 
pointto ACTIVE StoneMagw gs=40 PA=0.
   Years to correct for proper motions: 9.3579908675799
   Offset: StoneMagw->GS40 = 136.67 arcsec (54.05 W 125.52 N)
   For PA = 0.00 -----------------------------
      PCS_XY coords:   32.430   75.312
      AGw_XY coords:  -32.430  537.188
   For PA = 0.00 -----------------------------
      PCS_XY coords:   32.436   75.314
      AGw_XY coords:  -32.436  537.186
   For PA = 336.70 -----------------------------
      PCS_XY coords:    0.000   82.002
      AGw_XY coords:    0.000  530.498
   irc ClearStars
   irc SetHotspot 0.0 0.0
   irc SetOffset 0.0 0.0 RADEC
   irc SetStars target.dat
   irc PresetTelescope 0. POSITION ACTIVE left
...Done! (success)
   to examine the log: page LOGs/pointto_101059.log
0 H 1.6
!ssh agwuser@agw-control ./getdat1 40
        x: 32.360000
        y: 537.237000
Taking IRTC sequence 0.8 50 
   Single image obtained: irtc.20090510.101200.fits
ls: No match.
   Cube sequence obtained (maybe): Exit
...Done! (success)
/Repository/irtc.20090510_00012_cube.fit*
Transtar is finished with star 40 
Sun May 10 03:12:51 MST 2009

03:14 cl < tran_StoneM_00.cl which contains the following hardwired IRAF commands where the arguments are the star number. You have to set the other parameters for transtar before running this script.
# AGW transformation star script for Stone M field
# These stars are for PA=0
transtar (39)
transtar (40)
transtar (51)
transtar (72)
transtar (68)       irtc.20090510.102249.fits cube 0017
transtar (58)       irtc.20090510.102445.fits cube 0018 (an example of jumpiness)
transtar (52)       irtc.20090510.102637.fits cube 0019 (an example of jumpiness)
transtar (48)       irtc.20090510.102830.fits cube 0020
transtar (44)
transtar (43)
transtar (38)
transtar (37)
transtar (35)      irtc.20090510.103752.fits cube 0025
transtar (26)      irtc.20090510.103946.fits cube 0026
transtar (20)
transtar (13)      irtc.20090510.104331.fits cube 0028
transtar (18)      irtc.20090510.104525.fits cube 0029
transtar (51)      irtc.20090510.104723.fits cube 0030

Seeing is averaging about 0.6 arcsec FWHM on the 2 sec guider exposures.

This automated collection of transformation data may be one of the great inventions since sliced bread, although I don't suppose it will have much mass-market appeal. Now I suppose I will have to finish the draft of the April monthly report.

Based on watching this data come in, I'm starting to think that we still have some of the HBS type small jumps at the 0.25 arcsec level. In bad seeing, this is completely washed over by the atmospheric motion, but in good seeing the star image is more hoppy than I would expect from a Kolmogorov atmosphere. Before you draw to firm a conclusion, remember that we have been applying all the active corrections to M1 since the start of the AGW calibration data. The AGW transformation data should have been taken in split mode to miminize hardpoint jumps.

03:42 Seeing degrades to 1.3 arcsec. Does the automatic script need a break? This bad seeing happened for the last three stars in the above list.

03:48 pointto StoneMagw Tel="ACQUIRE" PAM- gs=0 - Adjust IE=-58 CA=+7

epar transtar for the next field
PACKAGE = IRTC
   TASK = transtar

gstarID =                   40  pointto-Guide star ID from catalog?
(sourceI=            StoneMagw) pointto-Source name from catalog?
(PA     =                  90.) pointto-Position angle to set to?
(agwNum =                    1) 0 for AGW3, 1 for AGW4
(exptime=                  0.8) takeseq-Exposure time in seconds?
(Nimages=                   50) takeseq-Number of images to take in sequence?
(verbose=                   no) Verbose output to screen?
(mode   =                   ql)

03:53 cl < tran_StoneM_90.cl

LBTO@lbtmu05:91 % more tran_StoneM_90.cl
# AGW transformation star script for Stone M field
# These stars are for PA=90
transtar (55)    irtc.20090510.105430.fits cube 0031
transtar (53)    irtc.20090510.105621.fits cube 0032
transtar (62)    irtc.20090510.105823.fits cube 0033
transtar (64)    irtc.20090510.110024.fits cube 0034
transtar (69)    
transtar (51)
transtar (58)    irtc.20090510.110700.fits cube 0037
transtar (68)    
transtar (72)    irtc.20090510.111105.fits cube 0039
transtar (85)    irtc.20090510.111311.fits cube 0040
transtar (83)
transtar (76)    irtc.20090510.111721.fits cube 0042
transtar (79)    irtc.20090510.111928.fits cube 0043
transtar (67) 
transtar (78)    irtc.20090510.112341.fits cube 0045
transtar (74)    irtc.20090510.112546.fits cube 0046
transtar (66)    irtc.20090510.112752.fits cube 0047
transtar (60)    irtc.20090510.112953.fits cube 0048
transtar (63)
transtar (61)    irtc.20090510.113404.fits cube 0050
transtar (54)    irtc.20090510.113602.fits cube 0051
transtar (49)    irtc.20090510.113807.fits cube 0052
transtar (59)    irtc.20090510.114011.fits cube 0053
transtar (50)    
transtar (55)

03:56 Put PSF active optics corrections back in "split" mode to minimize M1 motions starting with star 53.

03:58 Seeing is in the neighborhood of 1 arcsec, so this data isn't as good as the first data with PA=0 above.

04:46 offset 30 30 to move star off IRTC.

04:47 takeseq 0.8 50 for blank sky. irtc.20090510.114716.fits cube 0056

Close

04:50 Try to close for impending daytime. T=+8.7 degC D=-10.8 degC. Wind 5-8 m/sec.

04:51 TO-station freezes up. Wayne's screens are frozen. John's ssh window from lbtmu05 is unresponsive. lbtmu01 is not even pingable from lbtmu05. Reboot to-station the hard way (pressing the power button).

04:54 In the meantime, John launches an ECSGUI from lbtmu107 and starts closing the shutter doors.

05:00 Doors are closed, and Wayne is operating on the to-station again.

Short list of tasks for tomorrow night and following

  • Measure AGW- IRTC focus
  • More Transformation Data to test refined transformation
  • Offset tests with refined transformation
  • Check individual Zernike Calibrations (done the following night)
  • Open and Closed-loop tracking data
  • Try to get IRTC cubes of dome seeing
  • Vibration measurements on bright stars. (some done the following night)
  • Guiding tests versus magnitude
  • Check DG rotator scale
  • Collect better photometric data

The next version of this list is in Com20090513.

-- JohnHill - 10 May 2009
Topic revision: r12 - 11 Dec 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback